OpenStudy (waheguru):

CD Sell-Off EQME21282 Juan belongs to a CD club that sells CDs for $11.44 each before tax. His first shipment of CDs costs $90.49 including 13% tax. How many CDs are in his first shipment? Show your work

OpenStudy (waheguru):

i got 6.8

OpenStudy (cwrw238):

cost including tax = 11.44 * 1.13 = 12.9272 dollars 90.49 / 12.9272 = 7

OpenStudy (cwrw238):

to be exact its 6.999969 - thats pretty close to 7!